# Minimum Weighted Feedback Arc Sets for Ranking from Pairwise Comparisons

### v2

@@ Abstract
The problem of deriving global rankings from pairwise comparisons has received significant attention in the literature. Recently, [He et al. ICML2022] introduced learning-based methods for this task, demonstrating improvements over previous approaches. However, their experimental evaluation did not explore the connections with the closely related Minimum Weighted Feedback Arc Set (MWFAS) problem. This paper shows that simple, learning-free combinatorial algorithms for MWFAS outperform significantly these learning-based methods, achieving both faster runtimes and better ranking quality on the metrics and benchmarks established in [He et al. ICML2022]. [^1] This is a preliminary report.
@@ 1 Introduction
The ranking problem and the Minimum Weighted Feedback Arc Set (MWFAS) problem are closely related; however, the relationship between them has not been extensively explored in the literature. Our review of existing studies on ranking shows that only a few papers, such as [16] and [9], briefly mention the feedback arc set problem and a polynomial-time approximation scheme (PTAS) for tournaments [24], without conducting a thorough analysis of the connection between ranking and MWFAS. To better understand this relationship, we present definitions for both problems:
@@
Definition: The Ranking Problem Input: A directed graph $G=(V,E)$, where $V$ is the set of nodes representing agents or items to be ranked, and $E$ is the set of directed edges. Each edge $(i,j)\in E$ has a weight $w_{ij}>0$, representing the strength of preference or the victory of node $i$ over node $j$ in a pairwise comparison. Output: A ranking $R:V\to\mathbb{N}$ that assigns a rank $R(i)$ to each node $i\in V$. The ranking should minimize the total inconsistency with the pairwise comparisons represented by $E$. Specifically, the goal is to minimize the number of violated edges (or minimize the total weight of violated edges) where $R(i)>R(j)$ despite an edge $(i,j)$ indicating a preference for $i$ over $j$.
@@
Definition: The Minimum Weighted Feedback Arc Set Problem Input: A directed graph $G=(V,E)$, where $V$ is the set of nodes, and $E$ is the set of directed edges. Each edge $(i,j)\in E$ has an associated weight $w_{ij}>0$, representing the cost of removing that edge. Output: A subset of edges $E^{\prime}\subseteq E$ such that removing all edges in $E^{\prime}$ results in a Directed Acyclic Graph (DAG). The goal is to minimize the total weight of the removed edges, i.e., $\sum_{(i,j)\in E^{\prime}}w_{ij}$, thereby ensuring that the graph becomes acyclic with the least possible cost.
@@
The relationship between the ranking problem and the MWFAS problem is rooted in the goal of resolving inconsistencies within a directed graph to derive a meaningful ordering of nodes. In the ranking problem, the objective is to generate a ranking that aligns as closely as possible with given pairwise comparisons, minimizing the inconsistencies where ranks contradict preference directions. In a similar vein, the MWFAS problem focuses on eliminating cycles in the graph by removing the minimum weight of edges, ultimately transforming the graph into a Directed Acyclic Graph (DAG).
@@
This acyclic structure can then be used to perform a topological sort, yielding a ranking that adheres to the remaining pairwise comparisons. Therefore, the ranking problem can be reformulated as an instance of the MWFAS problem, wherein eliminating cycles allows for a consistent ranking that reflects the original preferences as much as possible. It is worth noting that the Minimum Feedback Arc Set problem, even in its unweighted version, is NP-hard [23], highlighting the inherent computational difficulty in deriving consistent rankings.
@@ 1.1 Ranking Problem and Minimum Weighted Feedback Arc Set (MWFAS)
The ranking problem can be effectively modeled as an instance of the Minimum Weighted Feedback Arc Set (MWFAS) problem. An edge from vertex $i$ to vertex $j$ with weight $w$ signifies that agent $i$ is preferred over agent $j$ by $w$ units. When the comparison graph is already a Directed Acyclic Graph (DAG), we can simply apply a topological sort to generate the ranking, where each edge from vertex $i$ to vertex $j$ ensures that $i$ is ranked higher than $j$.
In practice, however, comparison graphs often contain cycles, leading to inconsistencies in the pairwise preferences. To address these inconsistencies, we aim to delete the smallest possible subset of edges (in terms of total weight) such that the remaining graph is acyclic. This transformation corresponds directly to solving the MWFAS problem, where the goal is to remove edges with minimal cost to obtain an acyclic graph, thereby enabling a consistent ranking.
The MWFAS problem, like its unweighted counterpart, is known to be computationally challenging (NP-hard). While the Minimum Feedback Arc Set (MFAS) problem has been extensively studied in its unweighted form[32][25][14], the weighted version has received relatively less attention in the literature. The following sections discuss notable works aimed at solving the ranking problem.
@@ 1.2.1 Previous Work About Minimum Weighted Feedback Arc Set
  • The Minimum Weighted Feedback Arc Set (MWFAS) problem has received significant attention, particularly regarding approximation algorithms. [23] introduced this problem and proved it is an NP-HARD problem. Early work by [12] provided an approximation algorithm for the MWFAS problem, followed by improvements by [11], that introduced a divide-and-conquer approach using spreading metrics. Although effective, these algorithms used the ellipsoid method for solving linear programming relaxations, making them computationally slow in practice.
@@
    Later, [10] proposed an $O(VE)$ heuristic based on the local ratio method [4]. This heuristic was shown to be efficient in practice, providing a $\lambda$-approximation, where $\lambda$ is the length of the longest cycle in the graph.
@@
    [3] presented an exact method for solving the Minimum Feedback Arc Set (MFAS) problem with a focus on sparse directed graphs. They proposed a lazy cycle enumeration approach to iteratively extend an incomplete cycle matrix until an optimal solution was found. Given the NP-hard nature of the problem, their exact algorithm has exponential runtime and does not perform efficiently for large graphs.
    [21] proposed an $O(VE^{4})$ algorithm for MWFAS. However, as the number of edges in the graph increases, the algorithm’s performance becomes slower.
@@
    [29] addressed the parameterized complexity of the feedback vertex and arc set problems in tournaments, focusing on both weighted and unweighted versions. It presents fixed parameter tractable (FPT) algorithms for these problems, including an $O((2.4143)^{k}n^{\omega})$ algorithm for the weighted feedback vertex set problem and an $O((2.415)^{k}n^{\omega})$ algorithm for the weighted feedback arc set problem where $\omega$ is exponent of the running time of the best matrix multiplication algorithm, and $k$ is the size of the feedback set. The paper also introduces algorithms for the parametric duals of these problems and explores their complexity in oriented directed graphs and general directed graphs.
@@
    [19] addressed the problem of finding the largest subset of edges in a directed graph such that the resulting subgraph is acyclic. The authors propose randomized and deterministic approximation algorithms that achieve bounds better than previously known, specifically focusing on graphs without two-cycles and improving performance guarantees for low-degree graphs. The work also introduces methods for handling two-cycles optimally and establishes tight bounds for large classes of graphs.
@@
    [31] explored the use of satisfiability-modulo theories (SMT) solvers to solve the Minimum Feedback Arc Set (MFAS) problem. It presents an SMT encoding for MFAS and highlights the use of incremental solving features, such as unsatisfiable cores. The results show that while the SMT-based approach has potential, it remains computationally infeasible for larger instances, with solving times that scale poorly. The authors also provide complexity bounds, demonstrating that the MFAS problem remains challenging even with advanced SMT techniques.
1.2.2 Previous Work About Ranking Problem
The ranking problem has also been studied extensively in the literature.[22] introduced a novel ranking method based on pairwise comparisons, aimed at maximizing the probability of full agreement between a ranking and all corresponding pairwise comparisons. This method was demonstrated through portfolio analysis and was compared with previous approaches, highlighting its potential for improved prediction accuracy and indicating areas for further research.[9] introduced Sync-Rank, a method that employs eigenvector and semidefinite programming synchronization to solve ranking, constrained ranking, and rank aggregation problems. The approach leverages spectral or SDP relaxations, followed by rounding procedures to recover rankings. Sync-Rank outperformed other algorithms in both synthetic and real-world datasets while supporting semi-supervised ranking and identifying consistent partial rankings. [8] addressed the problem of ranking players using incomplete pairwise comparison data within the Bradley-Terry-Luce model framework. It established the minimax rate for this ranking task with respect to Kendall’s tau distance, identifying a unique transition between exponential and polynomial rates based on the signal-to-noise ratio. The paper proposed a divide-and-conquer algorithm that achieved the minimax rate, demonstrating its theoretical optimality and computational efficiency.
@@ 1.2.2 Previous Work About Ranking Problem
[30] proposed a simple, robust, and optimal ranking algorithm based on the Copeland counting method for pairwise comparisons. The algorithm ranks items by counting pairwise comparison wins, which is computationally efficient and achieves substantial speed-ups compared to previous methods. Unlike other models such as the Bradley-Terry-Luce model, this method does not rely on parametric assumptions and is robust to different data types.
@@
[28] presented the Distinguishing Feature (DF) model, which effectively learns from pairwise comparisons, even in cases involving intransitive preferences. The DF-Learn algorithm, which is based on a Siamese neural network, outperformed standard models on synthetic and real-world datasets while capturing complex cyclic dependencies.
@@
[18] proposed a GNN-based pairwise surrogate model to accelerate evolutionary optimization by predicting the superiority of candidate solutions rather than evaluating their exact fitness values. This model significantly reduced the number of fitness evaluations while maintaining comparable optimization performance. It is designed to work with any comparison-based evolutionary algorithm, offering a computationally efficient alternative for handling expensive optimization problems.
@@
[13] introduced Bias-Aware Ranker from Pairwise Comparisons (BARP), which extends the Bradley-Terry model with bias parameters to correct evaluator biases and produce unbiased rankings. BARP outperformed baseline methods on synthetic and real-world datasets.
@@
[17] introduceD a novel pipeline called Graph Neural Re-Ranking (GNRR), which integrates Graph Neural Networks (GNNs) into the re-ranking phase of information retrieval systems. By constructing corpus graphs to model semantic and lexical document relationships, the GNRR approach captures cross-document interactions, improving ranking performance on popular metrics compared to traditional uni-variate baselines. Experimental results show consistent performance gains, especially in challenging datasets, highlighting the potential of GNNs to refine document rankings effectively. The most similar work to ours is [6], which focuses on vertex ordering in directed graph streams. They addressed the Feedback Arc Set problem and propose removing feedback arcs for conducting topological sorts. However, unlike our work, their focus is on unweighted tournament graphs without experimental validation, whereas we address weighted general graphs with extensive experiments.
@@
[1] addresseD the feedback arc set problem and rank aggregation primarily in tournaments, a restricted class of directed graphs where every pair of vertices is connected by a directed edge. In contrast, our work generalizes the problem to arbitrary directed graphs, significantly broadening its applicability. Furthermore, while their study focuses on theoretical approximation algorithms without implementation or empirical validation, we provide practical heuristics accompanied by implementations and extensive experiments, demonstrating the effectiveness of our approach on real-world datasets.
@@
[5] presented exact algorithms for solving several vertex ordering problems on graphs. The proposed algorithms operate in either $O^{*}(2^{n})$ time and space or $O^{*}(4^{n})$ time and polynomial space, leveraging dynamic programming and recursive techniques similar to the Held-Karp and Gurevich-Shelah methods. These algorithms generalize approaches for the Travelling Salesman Problem and can be applied to a variety of vertex ordering problems, including treewidth and pathwidth. The main contribution lies in providing algorithms with moderately exponential complexity for a broad class of graph ordering problems.
@@
[7] addressed the problem of fair rank aggregation by introducing algorithms that ensure fairness across diverse groups in ranked outputs. The authors propose exact and approximation algorithms for the Closest Fair Ranking and Fair Rank Aggregation problems under various fairness notions, leveraging metrics such as Kendall tau and Ulam distances to achieve efficient and principled solutions.
2 A novel framework for ranking problem based on pairwise comparisons using MWFAS

### v3

@@ Abstract
The problem of deriving global rankings from pairwise comparisons has received substantial attention in the literature. Recent work by He et al. (ICML 2022) introduced learning-based approaches to this task, reporting improvements over existing methods. However, their experimental evaluation did not examine the relationship between ranking from pairwise comparisons and the closely related Minimum Weighted Feedback Arc Set (MWFAS) problem. In this paper, we demonstrate that simple, learning-free combinatorial algorithms for MWFAS can substantially outperform these learning-based approaches, achieving faster runtimes and higher ranking quality on the metrics and benchmark datasets established by He et al. (ICML 2022). [^1] This is a preliminary report.
@@ 1 Introduction
The ranking problem and the Minimum Weighted Feedback Arc Set (MWFAS) problem are closely related; however, the nature of this relationship has not been extensively explored in the literature. Prior work on ranking indicates that only a few studies, such as [16] and [9], briefly reference the feedback arc set problem and the existence of a polynomial-time approximation scheme (PTAS) for tournaments [24], without providing a detailed investigation of the connection to MWFAS. To better contextualize this relationship, we first present formal definitions of both problems:
Definition 1 (The Ranking Problem).
Let $G=(V,E)$ be a directed graph, where $V$ denotes the set of items to be ranked and $E$ is the set of directed edges. Each edge $(i,j)\in E$ has a weight $w_{ij}>0$, representing the strength of preference for $i$ over $j$ in a pairwise comparison.
@@
A ranking is a mapping $R:V\rightarrow\mathbb{N}$ that assigns an integer rank $R(i)$ to each $i\in V$. The objective is to minimize the total inconsistency with the pairwise comparisons represented by $E$. In particular, we aim to minimize the number (or total weight) of edges for which $R(i)>R(j)$ despite $(i,j)$ indicating a preference for $i$ over $j$.
@@
Definition (Minimum Weighted Feedback Arc Set Problem). Let $G=(V,E)$ be a directed graph, where $V$ is the set of vertices and $E$ is the set of directed edges. Each edge $(i,j)\in E$ has an associated weight $w_{ij}>0$, representing the cost of removing that edge from the graph.
The objective is to find a subset of edges $E^{\prime}\subseteq E$ such that removing all edges in $E^{\prime}$ yields a directed acyclic graph (DAG). The problem seeks to minimize the total weight of the removed edges, that is,
$\sum_{(i,j)\in E^{\prime}}w_{ij},$
thereby producing an acyclic graph with the smallest possible removal cost.
@@
The relationship between the ranking problem and the MWFAS problem is rooted in their shared objective of resolving inconsistencies in directed graphs to obtain a meaningful ordering of vertices. In the ranking problem, the goal is to produce an ordering that is as consistent as possible with the given pairwise comparisons, minimizing instances where the assigned ranks contradict the observed preferences. Similarly, the MWFAS problem aims to eliminate inconsistencies by removing a minimum-weight set of edges that break all directed cycles, thereby transforming the graph into a directed acyclic graph (DAG) from which a consistent ordering can be derived.
@@
Once the graph is made acyclic, a topological sort can be applied to obtain a ranking that is consistent with the remaining pairwise comparisons. Consequently, the ranking problem can be viewed as an instance of the MWFAS problem, in which cycles are removed to enable a consistent ordering that reflects the original preferences as closely as possible. It is important to note that the Minimum Feedback Arc Set problem, even in its unweighted form, is NP-hard [23], underscoring the inherent computational difficulty associated with deriving consistent rankings in general directed graphs.
@@ 1.1 Ranking Problem and Minimum Weighted Feedback Arc Set (MWFAS)
The ranking problem can be naturally modeled as an instance of the Minimum Weighted Feedback Arc Set (MWFAS) problem. In this formulation, a directed edge from vertex $i$ to vertex $j$ with weight $w$ indicates that item $i$ is preferred over item $j$ by a magnitude of $w$. If the comparison graph is already a directed acyclic graph (DAG), a topological sort directly yields a ranking in which each edge $(i,j)$ guarantees that $i$ is placed ahead of $j$ in the resulting order.
In practice, comparison graphs often contain cycles, resulting in inconsistencies among the pairwise preferences. To resolve these inconsistencies, the objective is to remove a subset of edges with minimum total weight such that the remaining graph becomes acyclic. This task corresponds directly to the MWFAS problem, which seeks to eliminate a minimal-weight set of edges to obtain a directed acyclic graph (DAG), from which a consistent ranking can then be derived.
The MWFAS problem, like its unweighted counterpart, is known to be computationally challenging, being NP-hard. While the Minimum Feedback Arc Set (MFAS) problem has been extensively investigated in its unweighted form [32, 25, 14], the weighted variant has received comparatively less attention in the literature. The following sections review notable contributions related to the ranking problem.
@@ 1.2.1 Previous Work About Minimum Weighted Feedback Arc Set
  • The Minimum Weighted Feedback Arc Set (MWFAS) problem has received considerable attention, particularly in the context of approximation algorithms. Karp [23] introduced the problem and established its NP-hardness. Early work by Even et al. [12] provided an approximation algorithm for MWFAS, which was later improved by Even, Naor, and Rao [11] through a divide-and-conquer approach based on spreading metrics. Although effective, these algorithms rely on the ellipsoid method for solving linear programming relaxations, which makes them computationally expensive in practice.
@@
    Later, Chakrabarti and Imamura [10] proposed an $O(VE)$ heuristic based on the local ratio method [4]. This approach was shown to be practical and efficient, achieving a $\lambda$-approximation, where $\lambda$ denotes the length of the longest cycle in the graph.
@@
    Baharev et al. [3] presented an exact method for solving the Minimum Feedback Arc Set (MFAS) problem, with a particular focus on sparse directed graphs. Their approach uses lazy cycle enumeration to iteratively extend an incomplete cycle matrix until an optimal solution is obtained. Due to the NP-hard nature of the problem, the algorithm exhibits exponential runtime and does not scale efficiently to large graphs.
  • Hecht et al. [21] proposed an $O(VE^{4})$ algorithm for MWFAS. However, the runtime grows rapidly as the number of edges increases, making the method impractical for large or dense graphs.
@@
  • Raman and Saurabh [29] investigated the parameterized complexity of the feedback vertex and feedback arc set problems in tournaments, considering both weighted and unweighted variants. They presented fixed-parameter tractable (FPT) algorithms, including an $O((2.4143)^{k}n^{\omega})$ algorithm for the weighted feedback vertex set problem and an $O((2.415)^{k}n^{\omega})$ algorithm for the weighted feedback arc set problem, where $\omega$ is the exponent of matrix multiplication and $k$ is the size of the feedback set. The authors also introduced algorithms for the parametric duals of these problems and examined their complexity in oriented tournaments and general directed graphs.
@@
  • Hassin and Rubinstein [19] studied the problem of identifying the largest subset of edges in a directed graph such that the induced subgraph is acyclic. They proposed randomized and deterministic approximation algorithms that achieve improved bounds over previous results, particularly for graphs without two-cycles and for low-degree graphs. Their work also introduced techniques for handling two-cycles optimally and established tight bounds for large classes of directed graphs.
@@
  • Shea-Blymyer and Young [31] investigated the use of satisfiability-modulo-theories (SMT) solvers for solving the Minimum Feedback Arc Set (MFAS) problem. They developed an SMT encoding that exploits incremental solving features, such as unsatisfiable cores. Although promising, their approach remains computationally infeasible for large instances, as solving times scale poorly. The authors also established complexity bounds, demonstrating that MFAS remains challenging even with advanced SMT-based techniques.
1.2.2 Previous Work on the Ranking Problem
The ranking problem has also been extensively investigated in the literature. Hochberg and Rabinovitch [22] introduced a ranking method based on pairwise comparisons that aims to maximize the probability of full agreement between a ranking and all corresponding pairwise relationships. Their method was evaluated through portfolio analysis and compared with previous approaches, demonstrating improved prediction accuracy and identifying opportunities for further research.
Cucuringu [9] proposed Sync-Rank, a method that employs eigenvector and semidefinite programming (SDP) synchronization to address ranking, constrained ranking, and rank aggregation tasks. The approach leverages spectral or SDP relaxations, followed by rounding procedures to recover rankings. Sync-Rank outperformed several competing methods on both synthetic and real-world datasets, while also supporting semi-supervised ranking and identifying consistent partial rankings.
Chen, Gao, and Zhang [8] studied the ranking of items from incomplete pairwise comparison data within the Bradley–Terry–Luce model. They established the minimax rate for this task with respect to Kendall’s tau distance, identifying a phase transition between exponential and polynomial rates depending on the signal-to-noise ratio. The authors proposed a divide-and-conquer algorithm that achieves the minimax rate, demonstrating both theoretical optimality and computational efficiency.
@@ 1.2.2 Previous Work on the Ranking Problem
Shah and Wainwright [30] introduced a simple, robust, and statistically optimal ranking algorithm based on the Copeland counting method for pairwise comparisons. The algorithm ranks items by counting wins in the comparison graph, offering substantial computational speed-ups over previous approaches. Unlike parametric models such as Bradley–Terry–Luce, the Copeland method does not rely on distributional assumptions and is robust to heterogeneous data types.
@@
Park and Ahn [28] introduced the Distinguishing Feature (DF) model, which learns from pairwise comparisons even in the presence of intransitive preferences. Their DF-Learn algorithm, based on a Siamese neural network architecture, outperformed standard models on both synthetic and real-world datasets while capturing complex cyclic dependencies among items.
@@
Gong and Rasingh [18] proposed a graph neural network (GNN)–based pairwise surrogate model to accelerate evolutionary optimization by predicting the superiority of candidate solutions rather than directly evaluating their fitness values. Their approach substantially reduced the number of required fitness evaluations while maintaining comparable optimization performance, and is compatible with a wide range of comparison-based evolutionary algorithms, providing a computationally efficient alternative for expensive optimization tasks.
@@
Fang and Bazaza [13] introduced Bias-Aware Ranker from Pairwise Comparisons (BARP), an extension of the Bradley–Terry model that incorporates bias parameters to correct evaluator biases and produce unbiased rankings. BARP outperformed several baseline methods on both synthetic and real-world datasets.
@@
Feng and Guo [17] introduced Graph Neural Re-Ranking (GNRR), a pipeline that integrates graph neural networks (GNNs) into the re-ranking stage of information retrieval systems. By constructing corpus graphs to model semantic and lexical document relationships, GNRR captures cross-document interactions and improves ranking performance on standard evaluation metrics. Experimental results show consistent performance gains, particularly on challenging datasets, highlighting the potential of GNNs to refine document rankings effectively.
The work most closely related to ours is by Chen and Ghosh [6], who studied vertex ordering in directed graph streams. They addressed the Feedback Arc Set problem by removing feedback arcs to enable topological sorting. However, their focus is on unweighted tournament graphs and lacks empirical validation, whereas we address weighted general graphs and provide extensive experimental evaluation.
@@
Alon [1] addressed the feedback arc set problem and rank aggregation primarily in tournaments, a restricted class of directed graphs in which every pair of vertices is connected by a directed edge. In contrast, our work generalizes the problem to arbitrary directed graphs, significantly broadening its applicability. Moreover, while Alon’s study focuses on theoretical approximation algorithms without implementation or empirical validation, we provide practical heuristics accompanied by implementations and extensive experiments, demonstrating the effectiveness of our approach on real-world datasets.
@@
Bodlaender and Fomin [5] presented exact algorithms for several vertex ordering problems on graphs. Their methods run in either $O^{*}(2^{n})$ time and space or $O^{*}(4^{n})$ time and polynomial space, leveraging dynamic programming and recursive techniques akin to the Held–Karp and Gurevich–Shelah approaches. These algorithms generalize techniques used for the Traveling Salesman Problem and can be applied to a variety of vertex ordering tasks, including treewidth and pathwidth. The key contribution of their work is the development of algorithms with moderately exponential complexity applicable to a broad class of graph ordering problems.
@@
Chen [7] studied the problem of fair rank aggregation by introducing algorithms that enforce fairness across diverse groups in ranked outputs. The author proposed both exact and approximation algorithms for the Closest Fair Ranking and Fair Rank Aggregation problems under several fairness notions, leveraging distance measures such as Kendall’s tau and Ulam distance to obtain efficient and principled solutions.
2 A Framework for Ranking from Pairwise Comparisons via MWFAS